OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x + y = 15 4x + 3y = 38

OpenStudy (anonymous):

x = 15-y 4(15-y)+3y=38 60-4y+3y =38 60-y=38 60-38 = y 22 = y x+22 = 15 x= 15-22 x=-7

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

x + y = 15 x +y-y = 15-y x = 15-y 4x+3y=38 4(15-y)+3y=38 60-4y+3y=38 60-y=38 60-y-38=0 60-y-38+y=y 22 = y x+y = 15 x+22=15 x+22-22 = 15-22 x = 7

OpenStudy (anonymous):

is the (7,22) the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

15-22 = -7 sry

OpenStudy (anonymous):

x+y = 15 (Equation 1) 4x +3y = 38 (Equation 2) Rearange Equation 1 to make it equation to x x = 15 - y, we shall call this Equation 3. We Substitute Equation 3, into Equation 2. This means we the x values in Equation 2, with the values x is equal to in Equation 3. Hence, 4 (15 - y) + 3y = 38 Expanding we get, 60 - 4y + 3y = 38 Simplifying, -y = - 22 Making this positive you get y = 22 Now we have a value for y, we can simply subsitute this into another Equation, I reccomend Equation 3. y = 22 x= 15 -y Subsituting our value for y x = 15 -22 x = -7 In the form you ask, (-7, 22) Hope this helps :)